\newcommand{\noop}[1]{}
\ifx\undefined\bysame
\newcommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\,}
\fi
\begin{thebibliography}{}

\bibitem[Andreas et~al., 2016]{Andreas-et-al-2016}
Andreas, Jacob, Marcus Rohrbach, Trevor Darrell, and Dan Klein (2016).
\newblock {Neural Module Networks}.
\newblock In {\em Proceedings of 2016 IEEE Conference on Computer Vision and
  Pattern Recognition (CVPR)}.

\bibitem[Arnold and Zuberb{\" u}hler, 2006a]{Arnold-Zuberbuhler-2006a}
Arnold, Kate and Klaus Zuberb{\" u}hler (2006a).
\newblock {Language Evolution: Semantic Combinations in Primate Calls}.
\newblock {\em Nature}, 441:~303.

\bibitem[Arnold and Zuberb{\" u}hler, 2006b]{Arnold-Zuberbuhler-2006b}
Arnold, Kate and Klaus Zuberb{\" u}hler (2006b).
\newblock {The Alarm-Calling System of Adult Male Putty-Nosed Monkeys, {\it
  Cercopithecus nictitans martini}}.
\newblock {\em Animal Behaviour}, 72:~643--653.

\bibitem[Arnold and Zuberb{\" u}hler, 2008]{Arnold-Zuberbuhler-2008}
Arnold, Kate and Klaus Zuberb{\" u}hler (2008).
\newblock {Meaningful Call Combinations in a Non-Human Primate}.
\newblock {\em Current Biology}, 18:~R202--R203.

\bibitem[Arnold and Zuberb{\" u}hler, 2013]{Arnold-Zuberbuhler-2013}
Arnold, Kate and Klaus Zuberb{\" u}hler (2013).
\newblock {Female Putty-Nosed Monkeys Use Experimentally Altered Contextual
  Information to Disambiguate the Cause of Male Alarm Calls}.
\newblock {\em PLOS ONE}, 8(6):~e65660.

\bibitem[Bahdanau et~al., 2018]{Bahdanau-et-al-2018}
Bahdanau, Dzmitry, Shikhar Murty, Michael Noukhovitch, Thien~Huu Nguyen, Harm
  de~Vries, and Aaron Courville (2018).
\newblock Systematic generalization: What is required and can it be learned?
\newblock arXiv:1881.12889v1.

\bibitem[Barrett, 2006]{Barrett-2006}
Barrett, Jeffrey (2006).
\newblock {Numerical Simulations of the Lewis Signaling Game: Learning
  Strategies, Pooling Equilibria, and Evolution of Grammar}.
\newblock {\em Technical Report, Institute for Mathematical Behavioral
  Science}.

\bibitem[Barrett, 2007]{Barrett-2007}
Barrett, Jeffrey (2007).
\newblock {Dynamic Partitioning and the Conventionality of Kinds}.
\newblock {\em Philosophy of Science}, 74:~527--546.

\bibitem[Barrett, 2009]{Barrett-2009}
Barrett, Jeffrey (2009).
\newblock {The Evolution of Coding in Signaling Games}.
\newblock {\em Theory and Decision}, 67:~223--237.

\bibitem[Barrett, 2016]{Barrett-2016}
Barrett, Jeffrey (2016).
\newblock {On the Evolution of Truth}.
\newblock {\em Erkenntnis}, 81:~1323--1332.

\bibitem[Barrett, 2017]{Barrett-2017}
Barrett, Jeffrey (2017).
\newblock {Truth and Probability in Evolutionary Games}.
\newblock {\em Journal of Experimental and Theoretical Artificial
  Intelligence}, 29(1):~219--225.

\bibitem[Barrett and LaCroix, 2020]{Barrett-LaCroix-2019}
Barrett, Jeffrey~A. and Travis LaCroix (2020).
\newblock Epistemology and the structure of language.
\newblock {\it Erkenntnis} (Forthcoming).

\bibitem[Barrett et~al., 2018]{Barrett-et-al-2018}
Barrett, Jeffrey~A., Brian Skyrms, and Calvin Cochran (2018).
\newblock {Hierarchical Models for the Evolution of Compositional Language}.
\newblock Unpublished Manuscript. May, 2018. PDF File.

\bibitem[Barrett et~al., 2017]{Barrett-et-al-2017}
Barrett, Jeffrey~A., Brian Skyrms, and Aydin Mohseni (2017).
\newblock {Self-Assembling Networks}.
\newblock {\em British Journal for the Philosophy of Science}.
\newblock Forthcoming.

\bibitem[Beck, 2017]{Beck-2017}
Beck, Jacob (2017).
\newblock {Can Bootstrapping Explain Concept Learning?}
\newblock {\em Cognition}, 158:~110--121.

\bibitem[Berwick and Chomsky, 2011]{Berwick-Chomsky-2011}
Berwick, Robert and Noam Chomsky (2011).
\newblock {The Biolinguistic Program: The Current State of its Development}.
\newblock In Sciullo, A. M.~Di and C.~Boeckx, editors, {\em The Biolinguistic
  Enterprise: New Perspectives on the Evolution and Nature of the Human
  Language Faculty}, pages 19--41. Oxford University Press, Oxford.

\bibitem[Berwick, 1998]{Berwick-1998}
Berwick, Robert~C. (1998).
\newblock {Language Evolution and the Minimalist Program: The Origins of
  Syntax}.
\newblock In Hurford, James~R., Michael Studdert-Kennedy, and Chris Knight,
  editors, {\em Approaches to the Evolution of Language: Social and Cognitive
  Bases}, pages 320--340. Cambridge University Press, Cambridge.

\bibitem[Berwick and Chomsky, 2016]{Berwick-Chomsky-2016}
Berwick, Robert~C. and Noam Chomsky (2016).
\newblock {\em {Why Only Us: Language and Evolution}}.
\newblock The MIT Press, Cambridge, MA.

\bibitem[Berwick et~al., 2013a]{Berwick-et-al-2013}
Berwick, Robert~C., Angela~D. Friederici, Noam Chomsky, and Johan~J. Bolhuis
  (2013a).
\newblock {Evolution, Brain, and the Nature of Language}.
\newblock {\em Trends in Cognitive Sciences}, 17(2):~89--98.

\bibitem[Berwick et~al., 2013b]{Berwick-Hauser-Tattersall-2013}
Berwick, Robert~C., Marc~D. Hauser, and Ian Tattersall (2013b).
\newblock {Neanderthal language? Just-So Stories Take Center Stage}.
\newblock {\em Frontiers in Psychology}, 4:~1--2.

\bibitem[Bickerton, 1990]{Bickerton-1990}
Bickerton, Derek (1990).
\newblock {\em {Language and Species}}.
\newblock University of Chicago Press, Chicago.

\bibitem[Bickerton, 1998]{Bickerton-1998}
Bickerton, Derek (1998).
\newblock {Catastrophic Evolution: The Case for a Single Step from
  Protolanguage to Full Human Language}.
\newblock In Hurford, James~R., Michael Studdert-Kennedy, and Chris Knight,
  editors, {\em Approaches to the Evolution of Language: Social and Cognitive
  Bases}, pages 341--58. Cambridge University Press, Cambridge.

\bibitem[Bolhuis et~al., 2014]{Bolhuis-et-al-2014}
Bolhuis, Johan~J., Ian Tattersall, Noam Chomsky, and Robert~C. Berwick (2014).
\newblock {How Could Language Have Evolved?}
\newblock {\em PLOS Biology}, 12(8):~e1001934.

\bibitem[Bush and Mosteller, 1955]{Bush-Mosteller-1955}
Bush, Robert~R. and Frederick Mosteller (1955).
\newblock {\em {Stochastic Models for Learning}}.
\newblock John Wiley \& Sons, Oxford.

\bibitem[Carey, 2004]{Carey-2004}
Carey, Susan (2004).
\newblock {Bootstrapping and the Origin of Concepts}.
\newblock {\em Daedalus}, 133(1):~59--68.

\bibitem[Carey, 2009a]{Carey-2009a}
Carey, Susan (2009a).
\newblock {\em {The Origin of Concepts}}.
\newblock Oxford Series in Cognitive Development. Oxford University Press,
  Oxford.

\bibitem[Carey, 2009b]{Carey-2009b}
Carey, Susan (2009b).
\newblock {Where our Number Concepts Come From}.
\newblock {\em The Journal of Philosophy}, 106(4):~220--254.

\bibitem[Carey, 2011a]{Carey-2011b}
Carey, Susan (2011a).
\newblock {Author's Response: Concept Innateness, Concept Continuity, and
  Bootstrapping}.
\newblock {\em Behavioral and Brain Sciences}, 34(3):~152--162.

\bibitem[Carey, 2011b]{Carey-2011a}
Carey, Susan (2011b).
\newblock {Precis of the Origin of Concepts}.
\newblock {\em Behavioral and Brain Sciences}, 34(3):~113--124.

\bibitem[Carey, 2014]{Carey-2014}
Carey, Susan (2014).
\newblock {On Learning New Primitives in the Language of Thought: Reply to
  Rey}.
\newblock {\em Mind \& Language}, 29(2):~133--166.

\bibitem[Carstairs-McCarthy, 1999]{Carstairs-McCarthy-1999}
Carstairs-McCarthy, Andrew (1999).
\newblock {\em {The Origins of Complex Language}}.
\newblock Oxford University Press, Oxford.

\bibitem[Choi et~al., 2018]{Choi-et-al-2018}
Choi, Edward, Angeliki Lazaridou, and Nando de~Freitas (2018).
\newblock Compositional obverter communication learning from raw visual input.
\newblock {\em arXiv preprint arXiv:1804.02341}.

\bibitem[Chomsky, 2002]{Chomsky-2002}
Chomsky, Noam (2002).
\newblock {\em On Nature and Language}.
\newblock Cambridge University Press, Cambridge.

\bibitem[Chomsky, 2005]{Chomsky-2005}
Chomsky, Noam (2005).
\newblock {Three Factors in Language Design}.
\newblock {\em Linguistic Inquiry}, 36:~1–22.

\bibitem[Chomsky, 2010]{Chomsky-2010}
Chomsky, Noam (2010).
\newblock {Some Simple Evo Devo Theses: How True Might They Be for Langauge?}
\newblock In Larson, Richard~K., Viviane D{\' e}prez, and Hiroko Yamakido,
  editors, {\em The Evolution of Human Language: Biolinguistic Perspectives},
  pages 148--162. Cambridge University Press, Cambridge.

\bibitem[Christiansen and Kirby, 2003]{Christiansen-Kirby-2003}
Christiansen, Morten~H. and Simon Kirby (2003).
\newblock {Language Evolution: The Hardest Problem in Science?}
\newblock In Christiansen, Morten~H. and Simon Kirby, editors, {\em Language
  Evolution}, pages 1--15. Oxford University Press, Oxford.

\bibitem[Culicover and Jackendoff, 2005]{Culicover-Jackendoff-2005}
Culicover, Peter~W. and Ray Jackendoff (2005).
\newblock {\em Simpler Syntax}.
\newblock Oxford University Press, New York.

\bibitem[Das et~al., 2017]{Das-et-al-2017}
Das, Abhishek, Satwik Kottur, Jos{\'e}~MF Moura, Stefan Lee, and Dhruv Batra
  (2017).
\newblock Learning cooperative visual dialog agents with deep reinforcement
  learning.
\newblock In {\em Proceedings of the IEEE International Conference on Computer
  Vision}, pages 2951--2960.

\bibitem[{Di Sciullo}, 2011]{Di-Sciullo-2011}
{Di Sciullo}, Anna~Maria (2011).
\newblock {A Biolinguistic Approach to Morphological Variation}.
\newblock In Sciullo, A. M.~Di and C.~Boeckx, editors, {\em The Biolinguistic
  Enterprise: New Perspectives on the Evolution and Nature of the Human
  Language Faculty}, pages 305--326. Oxford University Press, Oxford.

\bibitem[{Di Sciullo}, 2013]{Di-Sciullo-2013}
{Di Sciullo}, Anna~Maria (2013).
\newblock {Exocentric Compounds, Language and Proto-Language}.
\newblock {\em Language and Information Society}, 20:~1--26.

\bibitem[Erev and Roth, 1998]{Erev-Roth-1998}
Erev, Ido and Alvin~E. Roth (1998).
\newblock {Predicting How People Play Games: Reinforcement Learning in
  Experimental Games with Unique, Mixed Strategy Equilibria}.
\newblock {\em The American Economic Review}, 88(4):~848--881.

\bibitem[Evtimova et~al., 2017]{Evtimova-et-al-2017}
Evtimova, Katrina, Andrew Drozdov, Douwe Kiela, and Kyunghyun Cho (2017).
\newblock Emergent communication in a multi-modal, multi-step referential game.
\newblock {\em arXiv preprint arXiv:1705.10369}.

\bibitem[Fitch, 2004]{Fitch-2004}
Fitch, W.~Tecumseh (2004).
\newblock {Kin Selection and `Mother Tongues': A Neglected Component in
  Language Evolution}.
\newblock In Oller, D.~K. and U.~Griebel, editors, {\em Evolution of
  Communication Systems: A Comparative Approach}, pages 275--296. MIT Press,
  Cambridge, MA.

\bibitem[Fitch, 2010]{Fitch-2010}
Fitch, W.~Tecumseh (2010).
\newblock {\em {The Evolution of Language}}.
\newblock Cambridge University Press, Cambridge.

\bibitem[Fodor and Pylyshyn, 1988]{Fodor-Pylyshyn-1988}
Fodor, Jerry~A. and Zenon~W. Pylyshyn (1988).
\newblock {Connectionism and Cognitive Architecture: A Critical Analysis}.
\newblock {\em Cognition}, 28(1):~3--71.

\bibitem[Franke, 2014]{Franke-2014}
Franke, Michael (2014).
\newblock {Creative Compositionality from Reinforcement Learning in Signaling
  Games}.
\newblock In Cartmill, Erica~A., Se{\'e}n Roberts, Heidi Lyn, and Hannah
  Cornish, editors, {\em The Evolution of Language: Proceedings of the 10th
  International Conference}, volume~10 of {\em Evolang}, pages 82--89. World
  Scientific, Singapore.

\bibitem[Franke, 2016]{Franke-2016}
Franke, Michael (2016).
\newblock {The Evolution of Compositionality in Signaling Games}.
\newblock {\em Journal of Logic, Language and Information}, 25(3):~355--377.

\bibitem[Gaunt et~al., 2016]{Gaunt-et-al-2016}
Gaunt, Alexander~L., Marc Brockschmidt, Nate Kushman, and Daniel Tarlow (2016).
\newblock {Differentiable Programs with Neural Libraries}.
\newblock In {\em Proceedings of the 34th International Conference on Machine
  Learning}.
\newblock arXiv:1611.02109.

\bibitem[Giv{\' o}n, 1979]{Givon-1979}
Giv{\' o}n, Talmy (1979).
\newblock {\em {On Understanding Grammar}}.
\newblock Academic Press, New York.

\bibitem[Giv{\' o}n, 2002a]{Givon-2002a}
Giv{\' o}n, Talmy (2002a).
\newblock {\em {Bio-Linguistics: The Santa Barbara Lectures}}.
\newblock John Benjamins Publishing Company, Amsterdam.

\bibitem[Giv{\' o}n, 2002b]{Givon-2002b}
Giv{\' o}n, Talmy (2002b).
\newblock {The Visual Information-Processing System as an Evolutionary
  Precursor to Human Language}.
\newblock In Giv{\' o}n, Talmy and Bertram~F. Malle, editors, {\em The
  Evolution of Language Out of Pre-Language}, volume~53 of {\em Typological
  Studies in Language}, pages 3--50. John Benjamins, Amsterdam.

\bibitem[Giv{\' o}n, 2009]{Givon-2009}
Giv{\' o}n, Talmy (2009).
\newblock {\em {The Genesis of Syntactic Complexity: Diachrony, Ontogeny,
  Neuro-cognition, Evolution}}.
\newblock John Benjamins Publishing Company, Amsterdam/Philadelphia.

\bibitem[Hauser et~al., 2002]{Hauser-et-al-2002}
Hauser, Marc~D., Noam Chomsky, and W.~Tecumseh Fitch (2002).
\newblock {The Faculty of Language: What Is It, Who Has It, and How Did It
  Evolve?}
\newblock {\em Science}, 298:~1569--1579.

\bibitem[Havrylov and Titov, 2017]{Havrylov-Titov-2017}
Havrylov, Serhii and Ivan Titov (2017).
\newblock Emergence of language with multi-agent games: Learning to communicate
  with sequences of symbols.
\newblock In {\em Advances in neural information processing systems}, pages
  2149--2159.

\bibitem[Heine and Kuteva, 2007]{Heine-Kuteva-2007}
Heine, Bernd and Tania Kuteva (2007).
\newblock {\em {The Genesis of Grammar: A Reconstruction}}.
\newblock Oxford University Press, Oxford.

\bibitem[Hornstein, 2008]{Hornstein-2008}
Hornstein, Norbert (2008).
\newblock {\em {A Theory of Syntax: Minimal Operations and Universal Grammar}}.
\newblock Cambridge University Press, Cambridge.

\bibitem[Hu et~al., 2017]{Hu-et-al-2017}
Hu, Ronghang, Jacob Andreas, Marcus Rohrbach, Trevor Darrell, and Kate Saenko
  (2017).
\newblock {End-to-End Module Networks for Visual Question Answering}.
\newblock In {\em Proceedings of 2017 IEEE International Conference on Computer
  Vision}.
\newblock arXiv:1704.05526.

\bibitem[Hurford, 2007]{Hurford-2007}
Hurford, James~R. (2007).
\newblock {\em {Language in the Light of Evolution I: The Origins of Meaning}}.
\newblock Oxford University Press, Oxford.

\bibitem[Hurford, 2012]{Hurford-2012}
Hurford, James~R. (2012).
\newblock {\em {Language in the Light of Evolution II: The Origins of
  Grammar}}.
\newblock Oxford University Press, Oxford.

\bibitem[Huttegger, 2007]{Huttegger-2007b}
Huttegger, Simon~M. (2007).
\newblock {Evolutionary Explanations of Indicatives and Imperatives}.
\newblock {\em Erkenntnis}, 66:~409--436.

\bibitem[Jackendoff, 1999]{Jackendoff-1999}
Jackendoff, Ray~S. (1999).
\newblock {Possible Stages in the Evolution of the Language Capacity}.
\newblock {\em Trends in Cognitive Sciences}, 3:~272--279.

\bibitem[Jackendoff, 2002]{Jackendoff-2002}
Jackendoff, Ray~S. (2002).
\newblock {\em {Foundations of Language: Brain, Meaning, Grammar, Evolution}}.
\newblock Oxford University Press, Oxford.

\bibitem[Johnson et~al., 2016]{Johnson-et-al-2016}
Johnson, Justin, Bharath Hariharan, Laurens van~der Maaten, Li~Fei-Fei,
  C.~Lawrence Zitnick, and Ross Girshick (2016).
\newblock Clevr: A diagnostic dataset for compositional language and elementary
  visual reasoning.
\newblock In {\em Proceedings of 2017 IEEE Conference on Computer Vision and
  Pattern Recognition (CVPR)}.
\newblock arXiv:1612.06890.

\bibitem[Johnson et~al., 2017]{Johnson-et-al-2017}
Johnson, Justin, Bharath Hariharan, Laurens van~der Maaten, Judy Hoffman,
  Li~Fei-Fei, C.~Lawrence Zitnick, and Ross Girshick (2017).
\newblock Inferring and executing programs for visual reasoning.
\newblock In {\em Proceedings of 2017 IEEE International Conference on Computer
  Vision}.

\bibitem[Kamp and Partee, 1995]{Kamp-Partee-1995}
Kamp, Hans and Barbara Partee (1995).
\newblock {Prototype Theory and Compositionality}.
\newblock {\em Cognition}, 57:~129--191.

\bibitem[Kottur et~al., 2017]{Kottur-et-al-2017}
Kottur, Satwik, Jos{\'{e}} M.~F. Moura, Stefan Lee, and Dhruv Batra (2017).
\newblock Natural language does not emerge 'naturally' in multi-agent dialog.
\newblock {\em CoRR}, abs/1706.08502.

\bibitem[LaCroix, 2019a]{LaCroix-2019-Compositionality}
LaCroix, Travis (2019a).
\newblock Accounting for polysemy and role asymmetry in the evolution of
  compositional signals.
\newblock Unpublished Manuscript. May, 2019. PDF File.

\bibitem[LaCroix, 2019b]{LaCroix-2019-Signaling-Models}
LaCroix, Travis (2019b).
\newblock Evolutionary explanations of simple communication: Signalling games
  and their models.
\newblock {\em Journal for General Philosophy of Science / Zeitschrift f{\"u}r
  allgemeine Wissenschaftstheorie}.
\newblock Forthcoming.

\bibitem[LaCroix, 2019c]{LaCroix-If-Gradualism-2019}
LaCroix, Travis (2019c).
\newblock If gradualism is the correct approach ot language origins, then
  compositionality is not a plausible explanatory target.
\newblock Unpublished Manuscript. September, 2019. PDF File.

\bibitem[LaCroix, 2019d]{LaCroix-2019-Salt-v-Grad}
LaCroix, Travis (2019d).
\newblock Saltationist versus gradualist approaches to language origins: A
  critical discussion.
\newblock Unpublished Manuscript. March, 2019. PDF File.

\bibitem[LaCroix, 2019e]{LaCroix-2019-Logic-Game}
LaCroix, Travis (2019e).
\newblock Using logic to evolve more logic: Composing logical operators via
  self-assembly.
\newblock {\em British Journal for the Philosophy of Science}.
\newblock Forthcoming.

\bibitem[LaCroix, 2020]{LaCroix-Dissertation}
LaCroix, Travis (2020).
\newblock {\em Complex Signals: Reflexivity, Hierarchical Structure, and
  Modular Composition}.
\newblock PhD dissertation, University of California, Irvine.

\bibitem[Lake and Baroni, 2017]{Lake-Baroni-2017}
Lake, Brenden~M. and Marco Baroni (2017).
\newblock {Generalization Without Systematicity: On the Compositional Skills of
  Sequence-to-Sequence Recurrent Networks}.
\newblock arXiv:1711.00350.

\bibitem[Lazaridou et~al., 2018]{Lazaridou-et-al-2018}
Lazaridou, Angeliki, Karl~Moritz Hermann, Karl Tuyls, and Stephen Clark (2018).
\newblock Emergence of linguistic communication from referential games with
  symbolic and pixel input.
\newblock {\em CoRR}, abs/1804.03984.

\bibitem[Lazaridou et~al., 2016]{Lazaridou-et-al-2016}
Lazaridou, Angeliki, Alexander Peysakhovich, and Marco Baroni (2016).
\newblock Multi-agent cooperation and the emergence of (natural) language.
\newblock {\em arXiv preprint arXiv:1612.07182}.

\bibitem[Lewis, 1969]{Lewis-1969}
Lewis, David (2002/1969).
\newblock {\em {Convention: A Philosophical Study}}.
\newblock Blackwell, Oxford.

\bibitem[Lightfoot, 1991]{Lightfoot-1991}
Lightfoot, David (1991).
\newblock {Subjacency and Sex}.
\newblock {\em Language \& Communication}, 11(1--2):~67--69.

\bibitem[Margolis and Laurence, 2008]{Margolis-Laurence-2008}
Margolis, Eric and Stephen Laurence (2008).
\newblock {How to Learn the Natural Numbers: Inductive Inference and the
  Acquisition of Number Concepts}.
\newblock {\em Cognition}, 106(2):~924--939.

\bibitem[Margolis and Laurence, 2011]{Margolis-Laurence-2011}
Margolis, Eric and Stephen Laurence (2011).
\newblock {Beyond the Building Blocks Model}.
\newblock {\em Behavioral and Brain Sciences}, 34(3):~139--140.

\bibitem[Mart{\'i}nez, 2015]{Martinez-2015}
Mart{\'i}nez, Manolo (2015).
\newblock {Deception in Sender-Receiver Games}.
\newblock {\em Erkenntnis}, 80(1):~215--227.

\bibitem[Miyagawa, 2017]{Miyagawa-2017}
Miyagawa, Shigeru (2017).
\newblock {Integration Hypothesis: A Parallel Model of Language Development in
  Evolution}.
\newblock In Watanabe, S., M.~Hofman, and T.~Shimizu, editors, {\em Evolution
  of the Brain, Cognition, and Emotion in Vertebrates}, Brain Science, pages
  225--247. Springer, Tokyo.

\bibitem[Miyagawa et~al., 2015]{Miyagawa-et-al-2014}
Miyagawa, Shigeru, Shiro Ojima, Robert~C. Berwick, and Kazuo Okanoya (2015).
\newblock {The Integration Hypothesis of Human Language Evolution and the
  Nature of Contemporary Languages}.
\newblock {\em Frontiers in psychology}, 39:~564.

\bibitem[Moro, 2008]{Moro-2008}
Moro, Andrea (2008).
\newblock {\em {The Boundaries of Babel: The Brain and the Enigma of Impossible
  Languages}}.
\newblock The MIT Press, Cambridge, MA.

\bibitem[Neftci and Averbeck, 2019]{Neftci-Averbeck-2019}
Neftci, Emre~O. and Bruno~B. Averbeck (2019).
\newblock {Reinforcement Learning in Artificial and Biological Systems}.
\newblock {\em Nature Machine Intelligence}, 1:~133--143.

\bibitem[Newmeyer, 1991]{Newmeyer-1991}
Newmeyer, Frederick~J. (1991).
\newblock {Functional Explanation in Linguistics and the Origin of Language}.
\newblock {\em Language and Communication}, 11:~1–28.

\bibitem[Newmeyer, 1998]{Newmeyer-1998}
Newmeyer, Frederick~J. (1998).
\newblock {On the supposed `counterfunctionality' of Universal Grammar: Some
  evolutionary implications}.
\newblock In Hurford, James~R., Michael Studdert-Kennedy, and Chris Knight,
  editors, {\em Approaches to the Evolution of Language: Social and Cognitive
  Bases}, pages 305--319. Cambridge University Press, Cambridge.

\bibitem[Newmeyer, 2005]{Newmeyer-2005}
Newmeyer, Frederick~J. (2005).
\newblock {\em {Possible and Probable Languages: A Generative Perspective on
  Linguistic Typology}}.
\newblock Oxford University Press, Oxford.

\bibitem[Nowak and Krakauer, 1999]{Nowak-Krakauer-1999}
Nowak, Martin~A. and David~C. Krakauer (1999).
\newblock {The Evolution of Language}.
\newblock {\em Proceedings of the National Academy of Sciences},
  96:~8028--8033.

\bibitem[O'Connor, 2014]{OConnor-2014}
O'Connor, Cailin (2014).
\newblock {The Evolution of Vagueness}.
\newblock {\em Erkenntnis}, 79(4):~707--727.

\bibitem[Pemantle and Skyrms, 2004]{Pemantle-Skyrms-2004}
Pemantle, Robin and Brian Skyrms (2004).
\newblock {Network Formation by Reinforcement Learning: the Long and the Medium
  Run}.
\newblock {\em Mathematical Social Sciences}, 48:~315--327.

\bibitem[Piantadosi et~al., 2012]{Piantadosi-et-al-2012}
Piantadosi, Steven~T., Joshua~B. Tenenbaum, and Noah~D. Goodman (2012).
\newblock {Bootstrapping in a Language of Thought: A Formal Model of Numerical
  Concept Learning}.
\newblock {\em Cognition}, 123(2):~199--217.

\bibitem[Piattelli-Palmarini, 2010]{Piattelli-Palmarini-2010}
Piattelli-Palmarini, Massimo (2010).
\newblock {What Is Language, That It May Have Evolved, and What Is Evolution,
  That It May Apply to Language?}
\newblock In Larson, Richard~K., Viviane D{\' e}prez, and Hiroko Yamakido,
  editors, {\em The Evolution of Human Language: Biolinguistic Perspectives},
  pages 148--162. Cambridge University Press, Cambridge.

\bibitem[Piattelli-Palmarini and Uriagereka,
  2004]{Piattelli-Palmarini-Uriagereka-2004}
Piattelli-Palmarini, Massimo and Juan Uriagereka (2004).
\newblock {The Immune Syntax: The Evolution of the Language Virus}.
\newblock In Jenkins, Lyle, editor, {\em Variation and Universals in
  Biolinguistics}, volume~62 of {\em Linguistic Variations}, chapter~14, pages
  341--377. Elsevier, Oxford.

\bibitem[Piattelli-Palmarini and Uriagereka,
  2011]{Piattelli-Palmarini-Uriagereka-2011}
Piattelli-Palmarini, Massimo and Juan Uriagereka (2011).
\newblock {A Geneticist's Dream, a Linguist's Nightmare: The Case of FOXP2}.
\newblock In Sciullo, Anna Maria~Di and Cedric Boeckx, editors, {\em The
  Biolinguistic Enterprise: New Perspectives on the Evolution and Nature of the
  Human Language Faculty}, Oxford Studies in Biolinguistics, chapter~5, pages
  100--125. Oxford University Press, Oxford.

\bibitem[Pinker and Bloom, 1990]{Pinker-Bloom-1990}
Pinker, Steven and Paul Bloom (1990).
\newblock {Natural Language and Natural Selection}.
\newblock {\em Behavioral and Brain Sciences}, 13:~707--726.

\bibitem[Progovac, 2006]{Progovac-2006}
Progovac, Ljiljana (2006).
\newblock {The Syntax of Nonsententials: Small Clauses and Phrases at the
  Root}.
\newblock In Progovac, Ljiljana, Kate Paesani, Eugenia Casielles-Suarez, and
  Ellen Barton, editors, {\em The Syntax of Nonsententials: Multidisciplinary
  Perspectives}, volume~93 of {\em Linguistik Aktuell/Linguistics Today}, pages
  33--71. John Benjamins, Amsterdam.

\bibitem[Progovac, 2009a]{Progovac-2009a}
Progovac, Ljiljana (2009a).
\newblock {Layering of Grammar: Vestiges of Protosyntax in Present-Day
  Languages}.
\newblock In Sampsom, Geoffrey, David Gil, and Peter Trudgill, editors, {\em
  Language Complexity as an Evolving Variable}, Studies in the Evolution of
  Language. Oxford University Press, New York.

\bibitem[Progovac, 2009b]{Progovac-2009b}
Progovac, Ljiljana (2009b).
\newblock {Sex and Syntax: Subjacency Revisited}.
\newblock {\em Biolinguistics}, 3(2--3):~305–336.

\bibitem[Progovac, 2013]{Progovac-2013}
Progovac, Ljiljana (2013).
\newblock {Rigid Syntax, Rigid Sense: Absolutives/Unaccusatives as Evolutionary
  Precursors}.
\newblock In Franks, Steven, Markus Dickinson, George Fowler, Melissa Witcombe,
  and Ksenia Zanon, editors, {\em Proceedings of Formal Approaches to Slavic
  Linguistics (FASL), The Third Indiana Meeting, Bloomington, IN.}, pages
  246--259. Michigan Slavic Publications, Ann Arbor.

\bibitem[Progovac, 2015]{Progovac-2015}
Progovac, Ljiljana (2015).
\newblock {\em {Evolutionary Syntax}}.
\newblock Oxford University Press, Oxford.

\bibitem[Progovac, 2019]{Progovac-2019}
Progovac, Ljiljana (2019).
\newblock {\em {A Critical Introduction to Language Evolution: Current
  Controversies and Future Prospects}}.
\newblock Springer, Berlin.

\bibitem[Rescorla and Wagner, 1972]{Rescorla-Wagner-1972}
Rescorla, Robert~A. and Allan~R. Wagner (1972).
\newblock {A Theory of Pavlovian Conditioning: Variations in the Effectiveness
  of Reinforcement and Nonreinforcement}.
\newblock In Black, A.~H. and W.~F. Prokasy, editors, {\em Classical
  Conditioning II: Current Research and Theory}, pages 64--99.
  Appleton-Century-Crofts, New York.

\bibitem[Roth and Erev, 1995]{Roth-Erev-1995}
Roth, Alvin and Ido Erev (1995).
\newblock {Learning in Extensive Form Games: Experimental Data and Simple
  Dynamical Models in the Intermediate Term}.
\newblock {\em Games and Economic Behavior}, 8:~164--212.

\bibitem[Shea, 2009]{Shea-2009}
Shea, Nicholas (2009).
\newblock {New Concepts Can Be Learned}.
\newblock {\em Biology and Philosophy}, 26(1):~129--139.

\bibitem[Skyrms, 2010]{Skyrms-2010-Signals}
Skyrms, Brian (2010).
\newblock {\em {Signals: Evolution, Learning, \& Information}}.
\newblock Oxford University Press, Oxford.

\bibitem[Skyrms, 1996]{Skyrms-1996}
Skyrms, Brian (2014/1996).
\newblock {\em {Evolution of the Social Contract}}.
\newblock Cambridge University Press, Cambridge.

\bibitem[Skyrms and Barrett, 2018]{Skyrms-Barrett-2018}
Skyrms, Brian and Jeffrey~A. Barrett (2018).
\newblock {Propositional Content in Signals}.
\newblock {\em Studies in the History and Philosophy of Science C}.
\newblock Forthcoming.

\bibitem[Steinert-Threlkeld, 2014]{Steinert-Threlkeld-2014}
Steinert-Threlkeld, Shane (2014).
\newblock Learning to use function words in signaling games.
\newblock In Lorini, Emiliano and Laurent Perrussel, editors, {\em Proceedings
  of {\it Information Dynamics in Artificial Societies}}, (IDAS-14).

\bibitem[Steinert-Threlkeld, 2016]{Steinert-Threlkeld-2016}
Steinert-Threlkeld, Shane (2016).
\newblock Compositional signaling in a complex world.
\newblock {\em Journal of Logic, Language, and Information},
  25(3--4):~379--397.

\bibitem[Steinert-Threlkeld, 2020]{Steinert-Threlkeld-2018}
Steinert-Threlkeld, Shane (2020).
\newblock Towards the emergence of non-trivial compositionality.
\newblock {\it Philosophy of Science} (Forthcoming).

\bibitem[Szab{\' o}, 2012]{Szabo-2012}
Szab{\' o}, Zolt{\' a}n~Gendler (2012).
\newblock {The Case for Compositionality}.
\newblock In Hinzen, Wolfram, Edouard Machery, and Markus Werning, editors,
  {\em The Oxford Handbook of Compositionality}, pages 64--80. Oxford
  University Press, Oxford.

\bibitem[Sz{\'a}mad{\'o} and Szathmary, 2006]{Szamado-Szathmary-2006}
Sz{\'a}mad{\'o}, S. and E.~Szathmary (2006).
\newblock {Selective Scenarios for the Emergence of Natural Language}.
\newblock {\em Trends in Ecology \& Evolution}, 21(10):~555--561.

\bibitem[Tallerman, 2007]{Tallerman-2007}
Tallerman, Maggie (2007).
\newblock {Did our ancestors speak a holistic protolanguage?}
\newblock {\em Lingua}, 117(3):~579--604.

\bibitem[Tallerman, 2013a]{Tallerman-2013a}
Tallerman, Maggie (2013a).
\newblock {Join the Dots: A Musical Interlude in the Evolution of Language?}
\newblock {\em Journal of Linguistics}, 49:~455--487.

\bibitem[Tallerman, 2013b]{Tallerman-2013b}
Tallerman, Maggie (2013b).
\newblock {Kin Selection, Pedagogy, and Linguistic Complexity: Whence
  Protolanguage?}
\newblock In Botha, Rudolf and Martin Everaert, editors, {\em {The Evolutionary
  Emergence of Language}}, page 77–96. Oxford University Press, Oxford.

\bibitem[Tallerman, 2014a]{Tallerman-2014a}
Tallerman, Maggie (2014a).
\newblock {Is the Syntax Rubicon More of a Mirage? A Defence of Pre-Syntactic
  Protolanguage}.
\newblock In Cartmill, Erica~A., Se{\'e}n Roberts, Heidi Lyn, and Hannah
  Cornish, editors, {\em The Evolution of Language: Proceedings of the 10th
  International Conference}, volume~10 of {\em Evolang}, pages 318--325. World
  Scientific, Vienna.

\bibitem[Tallerman, 2014b]{Tallerman-2014b}
Tallerman, Maggie (2014b).
\newblock {The Evolutionary Origins of Syntax}.
\newblock In Carnie, Andrew, Yosuke Sato, and Daniel Siddiqi, editors, {\em
  {The Routledge Handbook of Syntax}}, pages 446--462. Routledge, London.

\bibitem[Tallerman and Gibson, 2011]{Tallerman-Gibson-2011}
Tallerman, M. and K.~Gibson (2011).
\newblock {\em {The Oxford Handbook of Language Evolution}}.
\newblock Oxford University Press, Oxford.

\bibitem[Wagner, 2014]{Wagner-2014}
Wagner, Elliott~O. (2014).
\newblock {Conventional Semantic Meaning in Signalling Games with Conflicting
  Interests}.
\newblock {\em British Journal for the Philosophy of Science}, 66(4):~751--773.

\bibitem[Wagner et~al., 2003]{Wagner-et-al-2003}
Wagner, Kyle, James~A. Reggia, Juan Uriagereka, and Gerald~S. Wilkinson (2003).
\newblock {Progress in the Simulation of Emergent Communication and Language}.
\newblock {\em Adaptive Behavior}, 11(1):~37--69.

\bibitem[Yang, 2013]{Yang-2013}
Yang, C. (2013).
\newblock {Ontogeny and phylogeny of language}.
\newblock {\em Proceedings of the National Academy of Sciences of the United
  States of America}, 110(16):~6323--6327.

\bibitem[Zollman, 2011]{Zollman-2011}
Zollman, Kevin J.~S. (2011).
\newblock {Separating Directives and Assertions Using Simple Signaling Games}.
\newblock {\em The Journal of Philosophy}, 108(3):~158--169.

\bibitem[Zollman et~al., 2012]{Zollman-et-al-2013}
Zollman, Kevin J.~S., Carl~T. Bergstrom, and Simon~M. Huttegger (2012).
\newblock {Between Cheap and Costly Signals: The Evolution of Partially Honest
  Communication}.
\newblock {\em Proceedings of the Royal Society B: Biological Sciences},
  280(1750):~20121878.

\bibitem[Zuberb{\"u}hler, 2001]{Zuberbuhler-2001}
Zuberb{\"u}hler, Klaus (2001).
\newblock {Predator-Specific Alarm Calls in Campbell's Monkeys, {\it
  Cercopithecus Campbelli}}.
\newblock {\em Behavioural Ecology and Sociobiology}, 50:~414--422.

\bibitem[Zuberb{\"u}hler, 2002]{Zuberbuhler-2002}
Zuberb{\"u}hler, Klaus (2002).
\newblock {A Syntactic Rule in Forest Monkey Communication}.
\newblock {\em Animal Behaviour}, 63(2):~293--299.

\end{thebibliography}

\documentclass{article}

     \usepackage[final]{neurips_2019}

\usepackage[utf8]{inputenc} % allow utf-8 input
\usepackage[T1]{fontenc}    % use 8-bit T1 fonts
\usepackage{hyperref}       % hyperlinks
\usepackage{url}            % simple URL typesetting
\usepackage{booktabs}       % professional-quality tables
\usepackage{amsfonts}       % blackboard math symbols
\usepackage{nicefrac}       % compact symbols for 1/2, etc.
\usepackage{microtype}      % microtypography
\usepackage{xcolor}
\usepackage{hyperref}
\hypersetup{
    colorlinks,
    linkcolor={red!50!black},
    citecolor={blue!50!black},
    urlcolor={blue!80!black}
}

\title{Biology and Compositionality: Empirical Considerations for Emergent-Communication Protocols}

\author{
  Travis~LaCroix\thanks{\href{travislacroix.github.io}{\texttt{http://travislacroix.github.io}}} \\
    \phantom{AND} \\
  Department of Logic \& Philosophy of Science\\
  University of California, Irvine\\
  Irvine, CA \\
  \phantom{AND} \\
  Mila \\
  Qu{\'e}bec AI Insitute \\
  Montr{\'e}al, PQ \\
  \phantom{AND} \\
  \texttt{tlacroix@uci.edu} \\
}

\begin{document}

\maketitle

\begin{abstract}
    Many researchers in language origins and emergent communication take {\it compositionality} as their primary target for explaining how simple communication systems can become more like natural language. % %Demonstrating how compositionality may have evolved is supposed to be (at least partially) sufficient for explaining how language evolved.
    I suggest that, if machine learning research in emergent communication is to take cues from biological systems and language origins, then compositionality is not the correct target.
\end{abstract}

\section{Introduction}

Communication is ubiquitous in nature, but {\it linguistic} communication---i.e., natural language---is unique to humans. There are inherent difficulties in determining how language may have evolved---i.e., out of simpler communicative precursors ({\it proto-language}). On the one hand, it is difficult to define what constitutes language---it is always in flux, it is infinitely flexible, and it is one of the most complex behaviours known \citep{Christiansen-Kirby-2003}. On the other hand, there is a paucity of direct evidence available with which to confirm current theories---language does not fossilise, nor can we go back in time to observe the actual precursors of human-level linguistic capacities.

Language origins research may appear to be a purely epistemic project; however, this programme finds application in contemporary machine learning and artificial intelligence (ML/AI). Understanding the fundamental principles that are involved in the (biological and cultural) evolution of effective communication may lead to innovative communication methods for use by interacting AI agents and multi-robot systems \citep{Wagner-et-al-2003}. AI agents need a common language to coordinate successfully and to communicate with humans. Furthermore, given the close relationship between language and thought, a clear understanding of how languages arise and persist in a population may also lend some additional insight into human cognitive capacities, helping lead the way in creating an artificial {\it general} intelligence (AGI).

The similar goals of language origins research (from a philosophical, biological, bio-linguistic, and cognitive systems perspective) and emergent communication research (from a computational perspective) should be apparent. One of the questions asked in contemporary machine learning is
\begin{center}
    {\it How can emergent communication become more like natural language?}
\end{center} 
This question finds parallel expression in language origins research:
\begin{center}
    {\it How might something like natural language evolve out of a simple communication system?}
\end{center}
Thus, progress in language origins may help direct research in emergent communication and vice-versa. However, there has been relatively little dialogue between these two approaches.

To explain how language evolved out of simpler precursors, it is necessary to understand the salient differences between language and simple communication, or signalling. Most researchers (in both programmes) working on these questions take {\it compositionality} (productivity, openness, hierarchical structure, generative capacity, complex syntax, etc.) as their primary target. Demonstrating how compositionality may have evolved is supposed to be (at least partially) sufficient for explaining how language evolved. Furthermore, narrowing one's explanatory target from the formidable question of how {\it language} evolved to the question of how one component of language---e.g.,  {\it complex syntax}---evolved affords some degree of analytic tractability. 

However, there is reason to think that compositionality is the wrong target, on the biological side (and so the wrong target on the ML side). As such, the purpose of this paper is to explore this claim. %, %
%the titular conditional statement, 
%focusing upon the consequences for potential research in emergent communication. %As a result, I will not argue about {\it whether} gradualism is the correct approach to language origins research; this will be beside the point since, of course, one cannot deny a conditional statement by denying the antecedent. Thus, I will assume that the antecedent is true and then proceed to discuss why the consequent follows from this assumption. 
This has theoretical implications for language origins research more generally, but the focus here will be the % significant 
implications for research on {\it emergent communication} in computer science and ML---specifically regarding the types of programmes that might be expected to work, and those which will not. My hope, then, is that this work will help to direct future research in fruitful new directions. 

\section{A Quick Note on Compositionality}

Compositionality provides a {\it prima facie} plausible explanatory target for language origins and emergent communication research. In the former case, this is because compositional syntax is a salient differentiating feature of language which is absent in simple communication systems found in nature. The {\it principle} of compositionality is typically stated as follows: {\it The meaning of a compound [complex] expression is a function of the meaning of its parts [constituents] and how they are combined [composed]} \citep{Kamp-Partee-1995, Szabo-2012}. In ML/AI, this is usually cashed out as a notion of {\it systematicity}; namely, ``the ability to entertain a given thought implies the ability to entertain thoughts with semantically related contents'' \citep[3]{Fodor-Pylyshyn-1988}. 

Any attempt to give an evolutionary explanation of human-level linguistic capacities will minimally need to account for the following: (1) how compositionality might arise from non-compositional communication; (2) if compositionality itself is an evolutionary adaptation, why compositional structure should be selected for in the first place; (3) why compositionality should be rare in nature, though communication is universal. %
Several models\footnote{See \citet{Nowak-Krakauer-1999, Barrett-2006, Barrett-2007, Barrett-2009, Franke-2014, Franke-2016, Steinert-Threlkeld-2014, Steinert-Threlkeld-2016, Steinert-Threlkeld-2018, Barrett-et-al-2018}.} have been suggested in recent years which grapple with these questions using the signalling-game framework.\footnote{The signalling game was introduced in \citet{Lewis-1969} and extended to dynamic models in \citet{Skyrms-1996, Skyrms-2010-Signals}. In the ML literature, this is sometimes called a {\it referential game}; see \citet{Lazaridou-et-al-2016, Das-et-al-2017, Evtimova-et-al-2017, Havrylov-Titov-2017, Kottur-et-al-2017, Choi-et-al-2018, Lazaridou-et-al-2018}.} Signalling games show how straightforward communication conventions might arise naturally through processes of repeated interactions. In an evolutionary context, starting with initially random signals and actions, individuals in a population learn or evolve effective communication conventions under several different dynamics. 

However, \citet{LaCroix-2019-Compositionality} suggests that the evolutionary explanations for compositional {\it signalling} offered thus far fail to give a plausible account of how compositionality might arise. The reason for this failure is twofold. On the one hand, these models often (if implicitly) take compositionality {\it qua} linguistic compositionality as their target for an evolutionary explanation. This gives rise to significant complications insofar as linguistic compositionality is rife with conceptual difficulties. On the other hand, these models fail to take into account the role asymmetry of the sender and receiver in the signalling game and thus fail to capture how compositionality might be beneficial for communication. \citet{LaCroix-2019-Compositionality} suggests that if compositionality is a good target of language origins (or emergent communication) research, then it is necessary to build a more straightforward notion of  compositionality for complex signalling from the bottom-up. Here, I suggest that compositionality is not even the right target for these research programmes. 

\section{Language Origins}

In the context of the salient distinctions between language and animal communication systems, the problem is that compositional signals are extremely rare, or nonexistent, in nature. There is scant empirical evidence for call sequences that are {\it compositionally} meaningful. Most current data that suggests compositional signalling in nature comes from Zuberb{\" u}hler's study of several species of African forest monkeys ({\it Cercopithecus}) \citep{Zuberbuhler-2001, Zuberbuhler-2002, Arnold-Zuberbuhler-2006a, Arnold-Zuberbuhler-2006b, Arnold-Zuberbuhler-2008, Arnold-Zuberbuhler-2013}. However, no such combinatorial capacity or anything similar has been found in {\it any} other species of monkey or great ape, though these are well-studied \citep{Fitch-2010}.

As a result, there is little evidence of a precursor to human-level compositional syntax. The most recent last common ancestor (LCA) between the great apes (including humans) and old-world monkeys (including {\it Cercopithecus}) existed approximately $25$ {\sc mya} (million years ago), during the Paleogene period. Since no other related species utilises such compositional capacities, it stands to reason that this syntactic disposition was {\it not} present in the LCA of {\it Homo} and {\it Cercopithecus}---the most recent common ancestor of humans and chimpanzees, for example, lived around $10-4$ {\sc mya}. Australopithecine species evolved after this break, with {\it Homo} likely evolving approximately $2$ {\sc mya}. 

Suppose that the LCA to great apes and old-world monkeys {\it did} have functionally proto-compositional syntax. This requires that proto-compositional syntax evolved up to $25$ {\sc mya}. Then, we should expect that most old-world monkeys {\it and} great apes would show signs of using proto-compositional syntax. However, as far as we can tell, they do not---excepting the few mentioned {\it Cercopithecus} species. Assuming these dispositions evolved in the LCA, $25$ {\sc mya}, and given that they do not appear in most decedents of the LCA, this implies that this disposition was lost---and lost more readily than it was held onto---in almost all other {\it Catarrhine} species.

Further still, it is controversial whether Neanderthals had language ($0.4-0.04$ {\sc mya}), let alone whether {\it H. heidelbergensis} had language ($0.7-0.2$ {\sc mya}). But, it is generally accepted that {\it Australopithecine} species ($3.9-2.9$ {\sc mya}) did {\it not} have language \citep{Fitch-2010}. This implies that this proto-compositional disposition evolved into fully-fledged natural language on the order of $0.2-2.0$ {\sc mya}.\footnote{\citet{Berwick-Chomsky-2016} give a more conservative estimate, between $0.06-0.2$ {\sc mya}, corresponding to the first anatomically-modern humans and the last exodus from Africa.} Meanwhile, these dispositions in few {\it Catarrhine} species remained utterly unchanged for more than $20$ million years, and these dispositions were lost in {\it almost all} great apes and old-world monkeys. This is the implausible (though technically not impossible) conclusion that follows from our assumption that a proto-compositional disposition existed in the LCA of great apes and old-world monkeys.  

Thus, the scant empirical evidence for compositional precursors to human language comes paired with an improbable evolutionary history. The more plausible alternative is that there is {\it no} empirical evidence for any precursor to human language in nature. Note that an account which posits the sudden emergence of compositional syntax does not fall prey to these criticisms since it requires no proto-compositional precursor to compositional language. This is a virtue of the so-called {\it saltationist} perspective on language origins. However, this view is not without its own problems \citep{LaCroix-2019-Salt-v-Grad}. Therefore, {\it if} gradualism is the correct approach to language origins, then compositionality {\it cannot} be a correct target.

\section{Consequences for Machine Learning}

Significant advances have been made in artificial systems by using biological systems as a guide. For example, the development of reinforcement learning algorithms was heavily inspired by learning rules that were studied empirically in biology \citep{Bush-Mosteller-1955, Rescorla-Wagner-1972, Roth-Erev-1995, Erev-Roth-1998}. A recent review of points of contact between ML and biology suggests several different areas of future research which can benefit from a bidirectional flow of information between these fields \citep{Neftci-Averbeck-2019}; however, nowhere is mentioned the interface between biological and computational models of the emergence of language.

I suggested above that a gradualist approach to language origins in biology implies that compositionality is an incorrect target for this research programme. Further, computer scientists working on emergent communication also take compositionality as a primary target---a benchmark for success. However, recent work in ML highlights several problems for learning compositional linguistic structures. In particular, neural networks latch on to statistical regularities in existing datasets: \citet{Bahdanau-et-al-2018} highlight that, in a synthetic instruction-following task \citep{Lake-Baroni-2017}, the agent does not learn a generalisation for composing words. Thus, when the agent is trained on the commands `jump', `run twice', and `walk twice', it subsequently fails when asked to interpret `jump twice'. Thus, they fail to {\it generalise}. %
%
%Neural networks are the `workhorse' of natural language comprehension and generation---\citet{Bahdanau-et-al-2018} highlight that neural networks play a significant role in machine translation \citep{Wu-et-al-2016} and text generation \citep{Kannan-et-al-2016} in addition to exhibiting state-of-the-art performance on several benchmarks, including {\it Recognising Textual Entailment}, \citep{Gong-et-al-2017}, {\it Visual Question Answering}, \citep{Jiang-et-al-2018}, and {\it Reading Comprehension} \citep{Wang-et-al-2018}. However, they 
%
The above analysis suggests a possible reason why they do so fail: the target is compositional communication. Thus, if ML is to take cues from biological systems, it seems that this is not the correct direction for research in emergent communication.

\section{Where to Go From Here: Reflexivity}

Communication is a unique evolutionary process in the following sense: once a group of individuals has learned some set of simple communication conventions, those learned behaviours may be used to influence future communicative behaviours, giving rise to a feedback loop. When faced with a novel context, an individual can always learn a brand new disposition. However, the individual may learn to take advantage of previously evolved dispositions. Indeed, they may learn to take advantage of pre-evolved {\it communicative} dispositions to thereby influence the evolution of future communicative dispositions. This is a notion of {\it reflexivity}. Reflexivity, unlike compositionality, is consistent with a gradualist approach to language origins. %
Once actors exhibit such complexity, at a small scale, it may lead to a feedback loop between communication and cognition that, over time, gives rise to the complexity that we see in natural languages. Thus, this evolutionary story depends inherently on a notion of {\it conceptual bootstrapping} \citep{Carey-2004, Carey-2009a, Carey-2009b, Carey-2011b, Carey-2011a, Carey-2014, Shea-2009, Margolis-Laurence-2008, Margolis-Laurence-2011, Piantadosi-et-al-2012, Beck-2017}.

To improve performance on generalisation, researchers in ML might add modularity and structure to their designs \citep{Andreas-et-al-2016, Gaunt-et-al-2016}. In the case of the Neural Module Network paradigm, a neural network is assembled from several {\it modules}, each of which is supposed to perform a particular subtask of the input processing. \citet{Bahdanau-et-al-2018} note that although this modular approach is intuitively appealing, ``widespread adoption has been hindered by the large amount of domain knowledge that is required to decide or predict how the modules should be created \ldots and how they should be connected \ldots based on a natural language utterance''. See also, \citet{Andreas-et-al-2016, Johnson-et-al-2016, Johnson-et-al-2017, Hu-et-al-2017}. %
Insofar as reflexivity is an apt target for the biological evolution of linguistic communication, it may too provide some insights for modelling emergent communication in an artificial system. See further discussion in \citet{LaCroix-If-Gradualism-2019, LaCroix-Dissertation}.

\section{Further Resources}

The gradualist view (though not necessarily via natural selection) is endorsed by, e.g., \citet{ %
    Givon-1979, 
        Givon-2002a, 
        Givon-2002b, 
        Givon-2009, 
    Pinker-Bloom-1990, 
    Newmeyer-1991, 
        Newmeyer-1998, 
        Newmeyer-2005, 
    Jackendoff-1999, 
        Jackendoff-2002, 
    Carstairs-McCarthy-1999, 
    Fitch-2004, 
        Fitch-2010, 
    Culicover-Jackendoff-2005, 
    Szamado-Szathmary-2006,
    Progovac-2006, 
        Progovac-2009a, 
        Progovac-2009b, 
        Progovac-2013, 
        Progovac-2015, 
        Progovac-2019, 
    Tallerman-2007, 
        Tallerman-2013a, 
        Tallerman-2013b, 
        Tallerman-2014a, 
        Tallerman-2014b,
    Heine-Kuteva-2007, 
    Hurford-2007, 
        Hurford-2012,
    Tallerman-Gibson-2011,
    Yang-2013}, among many others.
The saltationist view is endorsed by \citet{%
    Berwick-1998, 
        Berwick-et-al-2013, 
        Berwick-Hauser-Tattersall-2013, 
    Bickerton-1990, 
        Bickerton-1998, 
    Lightfoot-1991, 
    Hauser-et-al-2002, 
    Chomsky-2002, 
        Chomsky-2005, 
        Chomsky-2010, 
    Piattelli-Palmarini-Uriagereka-2004,
        Piattelli-Palmarini-Uriagereka-2011, 
    Moro-2008, 
    Hornstein-2008, 
    Piattelli-Palmarini-2010, 
    Berwick-Chomsky-2011, 
        Berwick-Chomsky-2016, 
    Di-Sciullo-2011, 
        Di-Sciullo-2013, 
    Bolhuis-et-al-2014, 
    Miyagawa-et-al-2014, 
        Miyagawa-2017}, etc.

%I have not argued here for the virtues of the former approach over the latter; though see \citet{LaCroix-2019-Salt-v-Grad}.

This is to say nothing of the signalling-game framework \citep{Lewis-1969, Skyrms-1996, Skyrms-2010-Signals} in evolutionary game theory, which has seen a number of significant advances in a variety of philosophically interesting domains. These include, e.g., the difference between indicatives and imperatives \citep{Huttegger-2007b, Zollman-2011}; signalling in social dilemmas \citep{Wagner-2014}; network formation \citep{Pemantle-Skyrms-2004, Barrett-et-al-2017}; deception \citep{Zollman-et-al-2013, Martinez-2015, Skyrms-Barrett-2018}; meta-linguistic notions of truth and probability \citep{Barrett-2016, Barrett-2017}; syntactic structure and compositionality \citep{Franke-2016, Steinert-Threlkeld-2016, Barrett-et-al-2018, LaCroix-2019-Logic-Game}; vagueness \citep{OConnor-2014}; and epistemic representations, such as how the structure of one’s language evolves to maintain sensitivity to the structure of the world \citep{Barrett-LaCroix-2019}. See \citet{LaCroix-2019-Signaling-Models} for an overview.

\small
\bibliographystyle{apalikelike}
\bibliography{Biblio}

\end{document}